Foreverangel:

Can someone check my answer please? Answer: -1 38 5 34

Foreverangel:

1 attachment
SmokeyBrown:

B+C would give [-1, 6 1, 4] Then A * (B+C) would be [3, 2 * [-1, 6 -1, 4] 1, 4] Top-left corner would be (3*-1) + (2*1) = -3+2 = -1 Top-right corner would be (3*6)+(2*4) = 18+8 = 26 Bottom-left corner would be (-1*-1)+(4*1) = 1+4 = 5 Bottom-right corner would be (-1*6)+(4*4) = -6+16 = 10 So, I think the final result would be [-1, 26 5, 10] But it has been a while since I've done Matrix Multiplication, so I am prepared to be corrected, if necessary

Foreverangel:

Hmm for the first part wouldn't it be 7+3 which is 10?

Foreverangel:

Since we are adding not subtracting

SmokeyBrown:

@foreverangel wrote:
Hmm for the first part wouldn't it be 7+3 which is 10?
Oh, you are right! I don't know why I saw a negative sign that wasn't there! In that case... [3, 2 * [-1, 6 -1, 4] 1, 10] And... Yes, you'd be correct in your calculations lol

Foreverangel:

Thank you so much!

SmokeyBrown:

No problem at all :)